40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 21 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  1607   Tue Nov 10 12:00:59 2015 AntonioDailyProgressBEATPLL and EAOM noise generated

The north cavity is now in place with 48% visibility! 

Quote:

The intensity noise that I have noticed after the EAOM installation disappeard with the use of a PBS before it.

Has to be mentioned that the light coming from the EAOM becomes eliptical with following power ratios at the PBS after theEAOM:

Pinc = 1.59mW (p-polarized);

Ptrans = 1.30mW (p);

Prefl = 0.25mW (s);

 

We were steering the beam transmitted from the cavities with use of lenses. I have realigned (not easy thing with such space constrains).

Intensity noise is way better, although the overall noise did not improved as the intensity noise is not the limiting noise. I think we need to suppress

more free-running noise. Actually I will go back to our FSS servo because I think we need different performances (not for now).

 

I have accidentally misaligned the north cavity, better not to say how :-);

 

  1606   Tue Nov 10 09:29:10 2015 AntonioDailyProgressBEATPLL and EAOM noise generated

The intensity noise that I have noticed after the EAOM installation disappeard with the use of a PBS before it.

Has to be mentioned that the light coming from the EAOM becomes eliptical with following power ratios at the PBS after theEAOM:

Pinc = 1.59mW (p-polarized);

Ptrans = 1.30mW (p);

Prefl = 0.25mW (s);

 

We were steering the beam transmitted from the cavities with use of lenses. I have realigned (not easy thing with such space constrains).

Intensity noise is way better, although the overall noise did not improved as the intensity noise is not the limiting noise. I think we need to suppress

more free-running noise. Actually I will go back to our FSS servo because I think we need different performances (not for now).

 

I have accidentally misaligned the north cavity, better not to say how :-);

  1605   Wed Nov 4 11:28:37 2015 AntonioDailyProgressBEATBeat noise did not improve with the ISSs ON

Summary

========

I have measured the noise at the beat note with both the ISS servo activated —>

NO improvement compared to the past. It is actually a bit worse. However with the EAOM

in the South path I had to change sluggishly the PDH gain settings.

 

- I took two measurements with different Fdev at the Marconi (1kHz and 10kHz) and the noise

is the same mostly. It seems that the limitation at moment relies NOT in the Marconi noise.

 

Data

=====

controlfb2/home/data/20151104_PLL_noise

 

Attachment 1: PLL_noise_comparison_20151104.pdf
PLL_noise_comparison_20151104.pdf
  1604   Wed Nov 4 11:14:24 2015 AntonioDailyProgressISSISS on South path is "working"

Summary

========

The EAOM has been implemented again in the South path. I still see an intensity noise effect in transmission, but it is much less than what I have seen from the previous days.

However the loop is suppressing most of it (but I am not happy about this). The figure shows the intensity noise and its suppression (Vdc = ~80mV).

 

Setup

=====

The setup is the one that it is currently in use in the North path:

 

lambda/2 ———p————> AEOM —-—> lambda/4 ——> PBS —>lambda/2———s—>EOM

 

I have also tried a different setup with light entering in the EAOM at 45 degrees, but the loop does not show suppression of intensity noise. I do not explain why at moment.

 

Note: When the light entering the EAOM is p (the current setup) the light coming out of the modulator

is circular polarized. 15% of ‘p’ goes in ’s’. This is not happening in the north path where the light

remains mostly linear. I am not convinced that this EAOM is properly functioning.

 

 

Data

=====

controlfb2/home/data/20151104_ISS_South

 

 

Attachment 1: ISS_South_noise.pdf
ISS_South_noise.pdf
  1603   Wed Nov 4 10:54:30 2015 AntonioHowToDocumentationelog etiquette

Scusami :-(. D'ora in poi usero' solo PDF :-)!!!

Quote:

Mama mia! crying

Per favore, utilizzare solo PDF.

 

  1602   Sun Nov 1 20:50:36 2015 ranaHowToDocumentationelog etiquette

Mama mia! crying

Per favore, utilizzare solo PDF.

  1601   Sat Oct 31 21:24:19 2015 AntonioDailyProgressAOMEAOM implemented in the South path again (2)

The EAOM is again in place but not the lambda/4 after it yet.

 

I found out that the intensity noise issue that I found after the implementation of the EAOM

is due to the lambda/4 implemented after it. I did not figured out what it is happening there.

At moment with the only EAOM there is no intensity noise issue.

(As reminder the lambda/4 is supposed to be rotated to produce circular light, splitting the light

in s and p at the output of a PBS located after it)

 

However I have optimized the setup of the lambda/4 soon after the laser. The light coming 

out from the laser is optimized to be linear and the lambda/2 after it is rotated to produce 1mW

after the Faraday Isolator (for the moment). This setup will be changed with the use of PBSO to

dump the light.

  1600   Thu Oct 29 21:06:53 2015 AntonioDailyProgressLaserCurrent setup of the lasers (to be changed) and power measurements

Summary

While I was debugging the "high" intensity noise at the ISSPD north i have noticed some scattering from the FI (north laser). It seems relatively well aligned but I did not want to touch it for now.

However I have measured the power emitted by the north laser and it is 306mW. The current setup provides ~99.3% dumping of the light into it. It means that only 2mW is at the output of the FI

while all the rest is dumped in the way out ---> I want to change it as soon as possible.


TO BE DONE

The laser setup will be changed in a way that the lambda/4 will maximize the linearization of the light (whatever angle is) and lambda/2 will maximize the power in transmition at the FI. A lambda/2 and a PBS

will be placed either before or after the FI in order to send 2mW to the rest of the setup. Now the question is to take care of the type of PBS because the damage threshold can be too low:

 

I consider the following two options:

1. PBS with damage threshold of 100W/cm^2 @(532nm) --> The minimum radius of the beam at 306mW is r ~ 628um (taking care of a factor 2 of safe margin and a factor 2 for 1064nm)

2. PBSO with damage threshold of 1MW/cm^2 @(1064nm) --> The minimum radius at the same 306mW is r ~ 4um;

I do not know the size of the beam. I do not have the optics to measure it and at moment I am not sure about previous measurements.

 

I have measured the power of the North laser vs the power on the display.

 

 

Some settings:

Laser South

Power Display = 479mW; Measured P = 306mW

DC = 2.12A

ADj +2

T + 34.4965

Laser North

Power Display = 68mW; Measured P=230mW

DC = 2.08A

ADj 0

T + 26.465

  1599   Thu Oct 29 20:14:33 2015 AntonioDailyProgressAOMEAOM removed

In order to debug the intensity noise that I found after the installation of the EAOM in the South path I have removed it from the path. The ASD measured at the ISS photodiode

located in transmition of the South cavity is anyway higher than what we have in transmition at the north cavity. Tomorrow I will try to optimize the other two EOMs alignment located

in the South path and then implement the EAOM again. However I see a very high drifting in the beate note.

  1598   Mon Oct 26 17:24:59 2015 AidanSummaryRefCavThermal shield cap drawing - lower emissivity material

Here's the drawing for the shield. https://dcc.ligo.org/LIGO-D1500403

 

Quote:

https://dcc.ligo.org/DocDB/0120/D1500213/001/D1500213-v1.PDF

 

  1597   Mon Oct 26 13:47:18 2015 AidanSummaryRefCavThermal shield cap drawing - lower emissivity material

https://dcc.ligo.org/DocDB/0120/D1500213/001/D1500213-v1.PDF

  1596   Sun Oct 25 22:15:35 2015 AntonioSummaryEOMAEOM installed in the South path. Noise issue.

Summary

The AEOM has been installed in the South path replacing the EOM 21MHz used for the PMC. There is a high noise that I clearly see at the photodiode in transmission. 

 

When I have placed the AEOM in the path I have decided to take the alignment of the previous EOM as reference. Not ideal because the reference should be the incoming beam. The beam is not parallel to the table and it was decided to be as less as possible invasive. The mode matching and the alignment gave at that time 20% of visibility (at each polarization). After the installation parameters where unchanged. Later I have improved the alignment bringing the visibility at 30% for both the polarizations. After that, when everything was in place I have easily locked the cavity but the power in transmition was showing a very high noise. I have spent all the day trying to twick the alignment because and servo loop gain, but we need to solve this before going further. My back does not allow me to proceed for today.

 

NOTE:

I have also noted that the South Laser which is labeed 2W laser has the lambda/4 and the lambda/2 rotated in a way that at the output of FI we had few mm. I am not sure if damping the power at the FI is a good thing.

  1595   Sun Oct 25 21:52:43 2015 AntonioDailyProgressEOMBeam profile measurement

Before the installation of the AEOM in the South cavity I wanted to have look to the beam profile along the paths. EOMs provokes distortion of the beam shape which may affect our mode-matching. It is important to keep the beam very small (200-500um diameter).

I think they are ok in the North path, a bit less good for the south path. Anyway I am going to use the beam as it is for the AEOM in the South path, replacing the EOM 21MHz used for the PMC with the AEOM that will be used for the ISS.

 

The pictures show the beam profile with the measurement done and with some ABCD matrix simulation for North and South path. They should come with an optical layout which I will make as soon as I will get OMNIGRAFFLE. I use inkscape but I will avoid that in order to be compatible with Rana and Aidan.

Data:

fb2/home/controls/data/20151021_beam_profile

Attachment 2: Slide2.jpg
Slide2.jpg
  1594   Tue Oct 20 19:43:05 2015 Antonio/Aidan/EricDailyProgressopticOptics inventory

The work that we need to do for the TCN experiment is at the state where we cannot easily go further without improving the lab organization.

I have collected the optics (1064nm) and made an inventory which is on dokuwiki: https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:optics. At moment is little but I hope we can keep this updated and we can follow the policy that is written there.

The inventory is shown in the picture.

  • Yesterday and today some cleaning has been done by Eric and Aidan, removing big useless parts. The rack on the left is free now.
  1593   Sun Oct 18 20:15:55 2015 AntonioDailyProgressISSISS on the North path and new optical setup

Summary

========

 

The work done yesterday and today gave us a working ISS in the North path with a different optical setup.

———————————————————————————————————————————-

 

Yesterday I have made several attempts to make the ISS working again in the North path. I have failed until I have noticed

the following setup for the AEOM:

 

lambda/2 ———p————> AEOM —-—> lambda/4 ——> PBS —> lambda/4 ———s—>EOM…..

 

Because I was not in agreement and I have asked Evan and I have figured out that this setup was not what it was meant to be.

Today I have replaced the second lambda/4 with a lambda/2 which made the ISS North working. Now the setup is:

 

lambda/2 ———p————> AEOM —-—> lambda/4 ——> PBS —>lambda/2———s—>EOM

 

The spectrum shows the intensity noise measured at the photodiode in transmission of the North cavity. Not clear what is happening

with the dark noise.

 

Loop setup: PD North ---> SR560 --->AEOM

SR560 setup: Gain at 5e4, HP=30Hz; LP=1kHz;

 

From  Yesterday I have also measured:

 

1. TF from AEOM to PD North;

2. TF from PD North to PLL control signal (injecting noise on AEOM);

 

SOME of the THINGS that I think need TO BE DONE in a short term:

 

1. We should implement the second AEOM in the South path. My plan is to replace the EOM that currently is in the path with the

AEOM because the beam shape requirement are the same, so it will be faster, given that we will not use the PMC. 

 

2. We need to check the shape of the beams at the modulators (ALL) in order to figure out if the (A)EOM requirements are respected.

This could be important for beam deformations which affect the mode matching at the cavities.

 

3. Need of a working dataviewer/IRcamera and a beam profiler;

 

4. It is very important that we spend some time in organizing the lab. The amount of time spent for looking for things is becoming an

obstacle for a proper lab work.

 

5. We also need to consider about the height of the table. Aside note: After two days in the lab my back is very painful.

 

6. Implement two Faraday Isolators in order to use only one polarization; for the moment I am even tempted to use a 50/50 BS, j

ust for the moment.

 

Curiosity:

So far we have been locking the cavities on the resonance given by the S-polarized light.

 

Data

=====

controlfb2/data/2015101718_PLL_ISS/

 

  1592   Fri Oct 16 22:24:35 2015 AntonioSummaryBEATPLL noise and beat frequency at 50MHz

Summary

Today the beat frequency was very difficult to find. Something changed, and I am not sure yet how to drive

the frequency of the beat. However the beat frequency is at 50MHz while so far it was at 64MHz. The PLL lock

more difficult and noisier.

 

Note:

Sometimes today or Yesterday we have increased the temperature of the lab (~ 2 degrees F).

 

Data:

lab computer/control fb2/home/data/20151016_PLL_noise_50MHz

  1591   Thu Oct 15 10:09:16 2015 AidanSummaryComputersUpdated network diagram on Wiki

https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:resources:computing:network:menu

Removed a few old computers from the diagram (35W laser, old OPC server, PSL Workstation, HWS workstation). Added a few new ones (OPC server in TCS lab, New workstation in PSL lab, Acromag control box in PSL lab).

 

  1590   Wed Oct 14 20:30:27 2015 AntonioDailyProgressBEATPLL noise

Summary

========

 

New PLL noise measurements have been made with  Fdev=1KHz and Fdev=10kHz at the input range of the Marconi.

There is an improvement of a factor ~10 for frequencies above 200Hz. However the two set of measurement show

the same noise in Hz/sqrt(Hz) between them. The lock of the PLL has been done keeping the same gain; when Fdev = 10KHz

the sr560 gain is 50, and when Fdev = 1kHz the SR560 gain is 500. The latter setting provokes an increase of the noise in V/sqrt(Hz)

which compensate for the reduction in the input range when Fdev goes from 10kHz to 1kHz. It is not clear to me what is happening.

 

Additionally it is not clear what is the “extra” noise that we see in the following noise budget as the sum does not match with the measurement.

=============================================================================================================

 

 

 

Plots:

=====

1. The first plot shows the noise budget with the measurement taken in September 12 (2015). We see that the PLL noise is dominated by the

Marconi and photo thermal noise;

 

2. The second plot shows the noise budget with the PLL noise measured today with Fdev=1kHz; Here I do not understand

what is happening above ~200Hz; the total noise is off from what has been measured. I also should check the power on the photodiode and

see if it matches the level (2dB) I have used for the Marconi noise measurement.

 

3. The same as in point 2. but here we have Fdev=10kHz; here the total noise is closer to match the measurement,

 

4. The last plot is just the comparison between the PLL noise taken in the past September and the one measured today.

 

Please note:  ISS is off

 

Data

====

Data are stored in TCN lab computer: controlfb2/data/20151014_PLL_noise

  1589   Tue Oct 13 16:47:55 2015 Aidan, AntonioSummaryEnvironmentPEM monitor computer is ready to run - needs license

We've set up the PEM monitor computer to broadcast the temperature channels to EPICS.

  • Computer is a Windows machine at 10.0.1.35
  • Has OPC server installed to communicate with the Newport zCDR wireless unit
  • OPC server still needs a license. Currently terminates after 2 hours
  • OpcIocShell runs to create an EPICS IOC that interfaces with the OPC server and broadcasts the channels via EPICS.

We'll find a permanent home for the machine tomorrow.

  1588   Mon Oct 12 16:22:19 2015 AntonioDailyProgressVCOMarconi ( IFR 2023a) OLTF, noise and calibration

I add a PLL transfer function taken with the SR785: It shows a unity gain frequency of ~54kHz instead of ~27kHz.

This is due to the AG4395  50 Ohm impedence (+ the 50 Ohm impedence at the output of the SR560). A factor of 2 is missing.

 

Quote:

Summary

=========

Before starting to work on reducing the noise, we decided to revise all subsystems in order to make sure that we know and understand the current state of each of them.

I made some noise measurements of the Marconi with two different carrier and different frequency range, obtaining the same results of I.D. PSL 816, 828, 874. Nothing

new at moment, so I will be very short and allocate the proceedure adopted in dokuwiki soon (https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:menu).

I resume the following results:

  • The calibration of the Marconi (South) for convertion in Hz units is : 711Hz/V;
  • Noise coming from Marconi can be reduced by lowering the input range independently from the carrier;
  • PLL open loop TF measured;
  • Electronic noise from Marconi measured.

--------------------->

The way I have measured is the same as described in the above mentione ELOG I.Ds.

 

Electronic noise:

=======

I have tried to measure the electronic noise from the Marconi but I am around 6 dB off. I would like to understand why...

I have used the same procedure that will be used to measure the photodiode electronic noise by disconnecting the feedback

in the PLL loop, measuring the noise of the control signal (Vfb) when no carrier is activated and by using the open loop  TF = DGA (D Discriminator G gain A actuator):

 

Vfb = 20e-6;

%Vfb = sqrt(Vfb^2 + Vfb^2)

sr = 10^(6/20);

A = 71.1;      % Hz/V

G = 2e3;       %

D = DGA/A/G;  

noise = Vfb/D/G

 

Data:

=====

All data are in the lab computer: controls/home/data/20151010_PLL_TF and 20151011_Marconi noise

 

Attachment 2: Mar_noise_2.png
Mar_noise_2.png
  1587   Sun Oct 11 23:16:34 2015 AntonioDailyProgressVCOMarconi ( IFR 2023a) OLTF, noise and calibration

Summary

=========

Before starting to work on reducing the noise, we decided to revise all subsystems in order to make sure that we know and understand the current state of each of them.

I made some noise measurements of the Marconi with two different carrier and different frequency range, obtaining the same results of I.D. PSL 816, 828, 874. Nothing

new at moment, so I will be very short and allocate the proceedure adopted in dokuwiki soon (https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:menu).

I resume the following results:

  • The calibration of the Marconi (South) for convertion in Hz units is : 711Hz/V;
  • Noise coming from Marconi can be reduced by lowering the input range independently from the carrier;
  • PLL open loop TF measured;
  • Electronic noise from Marconi measured.

--------------------->

The way I have measured is the same as described in the above mentione ELOG I.Ds.

 

Electronic noise:

=======

I have tried to measure the electronic noise from the Marconi but I am around 6 dB off. I would like to understand why...

I have used the same procedure that will be used to measure the photodiode electronic noise by disconnecting the feedback

in the PLL loop, measuring the noise of the control signal (Vfb) when no carrier is activated and by using the open loop  TF = DGA (D Discriminator G gain A actuator):

 

Vfb = 20e-6;

%Vfb = sqrt(Vfb^2 + Vfb^2)

sr = 10^(6/20);

A = 71.1;      % Hz/V

G = 2e3;       %

D = DGA/A/G;  

noise = Vfb/D/G

 

Data:

=====

All data are in the lab computer: controls/home/data/20151010_PLL_TF and 20151011_Marconi noise

Attachment 1: PLL_tf_plus6.png
PLL_tf_plus6.png
Attachment 2: PLL_tf_plus6.png
PLL_tf_plus6.png
  1586   Thu Oct 8 22:48:36 2015 ranaSummaryNoiseBudgetorder the heat shields

If you can't install python, you can run it all in the Sage Math Cloud for free.

Also, please post the final design for the heat shields which you've sent out for fab.

  1585   Thu Oct 8 16:27:18 2015 AidanSummaryNoiseBudgetNoise Budget on MATLAB - RIN is possibly the reason locking sucks right now

After two straight days of trying to get iPython notebooks to run on my Mac and two different Ubuntu installations, I gave up in frustration and rewrote the simpler parts of Evan's noise budget code into MATLAB. The following Noise Budget (without any BeatNote measurement from the lab) represents an estimate of the current noise in the system.

Clearly, with the ISS OFF, we are hugely dominated by intensity noise. We will investigate if this is the reason that we can't lower the FM deviation on the Marconi below 300kHz/V.

 

Attachment 1: CTN_NoiseBudget.pdf
CTN_NoiseBudget.pdf
Attachment 2: NBCode.zip
  1584   Tue Oct 6 09:27:25 2015 Aidan, EvanNotesNoiseBudgetiPython Noise Budget located in 40m SVN

/trunk/docs/algaas_ctn/sourceFiles/

  1583   Tue Oct 6 09:11:48 2015 Aidan, AntonioDailyProgressPEMIt's alive! Newport temperature sensors resurrected

We finally got the temperature sensors broadcasting to EPICS channels again - well, in part anyway. There are a lot of configuration issues to work out (refresh rate, saving to frames, license for OPC server, battery monitors, data precision). But at least we can now see a temperature sensor channel in EPICS that corresponds to a live measurement. The configuration to get the data from the remote unit to EPICS is shown in the attached block diagram.

 

More details can be found here:

https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:menu

Attachment 1: ZCDR_PEM_OPC.pdf
ZCDR_PEM_OPC.pdf
  1582   Sun Sep 13 13:56:04 2015 AntonioDailyProgressElectronics EquipmentGain Knobs calibration on the North side

Summary

========

In order to have a better understanding of the gain associated to the Knobs on the interface PDH box, I took some measurements from TestInput to Out1Fast on the FSS field box

at different gain values for the common and fast knobs and measured the gain of the transfer function at DC. From fitting dB vs Knob counts

we get:

 

North:

Common Knobs dB/100Counts = 2.35dB

Fast Knob: dB/100Counts = 2.49dB

 

I took more measurements than I needed of course, I did it to check while I was taking measurement that some unwanted electronic effect was happening.

 

Data

Data are on the lab control/home/data/20150831_Knobs_calibration/

 

Attachment 1: CommonKnob.pdf
CommonKnob.pdf
Attachment 2: CommonKnob.pdf
CommonKnob.pdf
Attachment 3: FastKnob.pdf
FastKnob.pdf
  1581   Sun Sep 13 12:44:14 2015 AntonioDailyProgressFSSOLTF of the north loop (now with data)

Summary

=========

I needed to have some data of the current OLG of one of the two PDH loop in order to fully model the electronic (+ optic setup).

The measured transfer function of the north loop has a unity gain around 150kHz at maximum gain settings.

The description of the measurement is in ID 1575. However now I know that the TF from Comm out1 to Comm out2 has gain of 1, this means

that we do not need to add this part of the electronic path when we measure the OLG.

 

---> At moment I did not succed yet with modeling the OLG.

 

Figure

it shows the TF of the north loop PDH at different value of the gains.

 

 

Data

Data are on the lab control/home/data/20150910_NorthPDH_transfer_functions/

 

Attachment 1: NorthOLG.pdf
NorthOLG.pdf
  1580   Sun Sep 13 11:31:04 2015 AntonioDailyProgressBEATPLL noise comparison at different PDH gain settings

Summary

========

Measurement of the PLL noise have been taken at different PDH gains settings. Noise start to increase when the gain on the oh North loop

are below North: Common/Fast=500/300 and South: Common/Fast=420/420. Stays fairly the same with higher values.

 

The measurement are still in V/sqrt(Hz), but for now I use them only for comparison purpose.

 

Figure:

Noise comparison at different gains: In the legend are the "values" of the gains (the numbers on the knobs)

Current measurements in Vrms/sqrt(Hz)

Beat freq = ~ 65 MHz;

FM deviation = 400kHz

gain = 20;

 

Data

Data are on the lab control/home/data/20150912_PLL_noise/

 

Attachment 1: NoiseComparison_20150912_Vrms.pdf
NoiseComparison_20150912_Vrms.pdf
Attachment 2: NoiseComparison_20150912_Vrms.pdf
NoiseComparison_20150912_Vrms.pdf
  1579   Sun Sep 13 10:40:29 2015 AntonioDailyProgressBEATPLL measured noise

Summary

========

The goal of the TCN experiment is to measure the TC noise. This requires to lower down the noise level that we have at

the PLL output. I have decided to take noise measurement today in order to have a reference from the level we are starting with.

We should start to implement changes in order to lower the noise at the output and keep monitoring it.

----------------------------------------------------------------------------------------------------------------------------------------------------------------------------

 

Description

=========

I report a plot of the PLL noise in Vrms/sqrt(Hz) as it is not clear yet how to convert the units in Hz/sqrt(Hz). From older Elogs ID 889,

I see that there is available a calibration for the IFR 2023a, and this depends on the input range. I am not sure on what the input range is.

The measurement has been taken in dirrent frequency range in order to have a better resolution. For now, when we are at this level

of noise it is not worth it.

 

Note:

I have tried to lock the PLL with a FM deviation which is less than 300kHz but it was not possible. There is too much noise.

 

Figure 1:

Target noise and noise taken sometimes in the past by Evan’s (in Hz/sqrt(Hz)).

 

Figure 2:

Current measurements in Vrms/sqrt(Hz);

Beat freq = ~ 65 MHz;

FM deviation = 400kHz

gain = 20;

note: something happened after the first measurement, but at this point is not so important.

 

Data:

Data are on the lab control/home/data/20150912_PLL_noise

 

 

/

Attachment 1: NoiseEvan.pdf
NoiseEvan.pdf
Attachment 2: Noise_20150912_Vrms.pdf
Noise_20150912_Vrms.pdf
  1578   Wed Sep 2 22:13:47 2015 AntonioSummaryBEATPLL locking AGAIN

Description

As described in the elog entry n. 1577 we were not able to lock the PLL as has been described in elog n. 1570. I have started by playing with the two PDH gains of both loops (North and South) as this could have been causes of non tolerable noise in the PLL loop. I have also monitored the peaks described in entry n.1577 as we were suspicious for their preventing the PLL locking.

Conclusion

The PLL loop has been locked repeatedly after locking the cavities multiple times. This result has been achieved by setting the PLL gain on the SR570 at 20 (PLEASE NOTE I was not able to lock the PLL with any other gains settings). 

The peakes of entry n.1577 are not preventing us to lock the PLL.

 

Some settings:

North FSS interface: Common gain =700; Fast gain = 450; PID = 4.451V;

South FSS interface: Common gain = 850; Fast gain =  250; PID = 0.5527V;

Beat frequency = 69.6 MHz;

 

However these are not the only allowable settings for the gain, but the PDH loop gains are crucial for the PLL locking. Later I am going to give a quantitative analysis for our PDH loops in order to have them in a more stable and/or less noisy locking point.

  1577   Thu Aug 27 17:11:43 2015 Aidan, AntonioSummaryBEATTrouble locking - 140kHz notes are switching supplies?

We tried locking the PLL today but failed. Looking at the beat note on the network analyzer revealed some 70/140kHz harmonics. We thought this might be what is preventing us from locking.

Rana suggested that these are from the switching power supplies (which switch at 70kHz). If so, this may be a red-herring. It's possible we're not trying with high enough gain settings ...

The large peak that is offset by about 50kHz is from the Marconi ... we can make it move around by changing the Marconi frequency.

Attachment 1: network_analyzer_140kHz.JPG
network_analyzer_140kHz.JPG
  1576   Tue Aug 25 19:12:36 2015 Antonio, EricDailyProgressComputersGPIB installed

Today Eric provided his Python scripts (and installed them) needed to connect the SR785 and the AG4395A devices with our lab computer through the GPIB interface. With these scripts we are able to download measurement data that we take by using the two above mentioned devices, plot them and set the measurement settings directly from the computer. Mainly we need to use two of them, i.e. with following commands:

1. AGmeasure: AGmeasure --getdata -i 10.0.0.13

2. SRmeasure: SRmeasure --getdata -i 10.0.0.13

These scripts can run from any folders on the computer. These and some other features will be explained in the TCN wiki page, which I am going to write soon.

We now need to make the lab computer accessible from other computers. The SSH protocol is iinstalled, but the modem need to be configured. Less attaractive but a possible option is to have a svn folder on the lab machine.

 

  1575   Thu Aug 20 13:55:22 2015 AntonioSummaryFSSOLTF Transfer functions of the North loop

In order to have a better understanding of the FSS electronic boards we decided to take

open loop transfer functions (OLTF) of the FSS North cavity loop. In the same time we took 

the opportunity to verify that the broadband EOM is working. However this measurement will

be done again for a better quantitative analysis of the loop performances.

 

OLTF measured in the common path shows the shape due to the PZT and due to the EOM

paths as expected.The UGF has been pushed up to 170kHz. In the picture below we see the

traces corrisponding to the two different gain settings.

 

Connections:

  

Network Analyzer                           FSS

           RF          ----------->        EXC (Common)

           R            ----------->        OUT2 (Common)

           A            ----------->        OUT1 (Common)

 

These measurements have been done with the Switch Exc on located on FSS interface board

and with the following gain settings:

 

  Fine Common Fast Offset

Initial settings

580 312 240 506
High gain 950 950 950 506

 

 

 

OLTF measured in the FAST path have the connections applied at the Fast and the switch EXC on

the front panel turned on. This measurement has been taken with "high gain settings" of the previous

measurement.In the following pictures can be seen the previous OLTF too (green/red).

 

Connections:

 

Network Analyzer                           FSS

           RF          ----------->        EXC (Fast)   (front panel switch Exc ON)

           R            ----------->        OUT2 (Fast)

           A            ----------->        OUT1 (Fast)

 

 

PLEASE note these are not the total OLTF (relative to their path) because is still missing the transfer function

between OUT1 and OUT2 in both the paths Common and Fast.

 

I would like to model this loop, I need to figure out the best way to do it.

  1574   Wed Aug 19 13:53:45 2015 AidanSummaryDAQDAC calibration for Acromag slow units

Volts (requested) to Counts

C = a1*V_req + a2

a1 = 3003.5 counts/Volt

a2 = -20.8285 counts

Entered counts (in +ve and -ve amounts) and measured the resulting voltage across the output. I'm not sure how stable this calibration is.

Attachment 1: Photo_on_8-19-15_at_1.51.33_PM.jpg
Photo_on_8-19-15_at_1.51.33_PM.jpg
  1573   Wed Aug 19 10:34:33 2015 AidanSummaryDAQBinary IO control of TTFSS box working - can lock PDH remotely

I've worked out how to get the binary IO to work with the TTFSS box so that we can activate switches in that unit. It wasn't working in the setup yesterday because of physics. Actually - there is a 10K pulldown resistor in the Acromag unit that attaches the output to ground. The actual circuit looks like this:

 

VCC (5V) --- (4.99K) --- T1EN -----|-----DIO0 ----(6.2V if DOUT set to 1)---- (10K) ------ | GND

.........   TTFSS...................................| ..............................ACROMAG ...............................................|

T1EN is measured by the switch-chip (SN74HCT157D, chip U3 in D040423) to determine whether it should be open or closed. We need to bring T1EN below 0.8V to get the TTL logic to work.

If DOUT is set to 1, then DIO0 and T1EN become the excitation voltage, 6.2V, and the switch circuit reads high. If DOUT is set to 0, the excitation voltage is removed and we just end up with a voltage divider and around 3.33V at T1EN - which does not register as low.

We can get around this by adding a smaller resistor, say 810 Ohms, in parallel to the 10K, to lower the effective resistance of the pull-down resistor to 750 Ohms. The maximum current the Acromag unit will have to supply is 6.2V/750 Ohms = 8.4mA.

So that's what I did. Now, when I switch DOUT to 1, I see 6.2V at T1EN and when I switch DOUT to 0, I see 0.669V at T1EN. The TTFSS box registers these as two different states and I can lock and unlock the PDH loop from EPICS.

  1572   Mon Aug 17 17:02:41 2015 AidanSummaryComputersAcromag binary outputs working - still can't control TTFSS box though

I managed to figure out the modbusDrv configuration settings to get the binary output of the Acromag working. I've updated the Wiki page to reflect this. I've wired the XT1541 DAC, BIO Acromag unit to the T1EN and T2EN channels on the TTFSS box but I still can't get remote control of it yet for some reason. When the PDH loop is closed and I switch the TTFSS box to REMOTE, the loop stays closed regardless of what I do to the binary outputs in EPICS.

https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:resources:computing:acromag

  1571   Fri Aug 14 19:10:39 2015 AidanSummaryTempCtrlDropped lab temperature set point from 72F to 68F

At 7:07PM

 

  1570   Thu Aug 13 09:07:04 2015 Aidan, Rich, AntonioSummaryBEATPLL locked on beat note - TTFSS boxes were being used incorrectly previously

We investigated the way we were locking the PDH loops using the TTFSS boxes. Here's what we previously did:

  • With the loop open and the TTFSS interface set to LOCAL & TEST, we adjusted the temperature of the laser until we were close to TEM00 resonance.
  • Then we set the first switch from LOCAL to REMOTE. 
  • This locked the loop.

We realized, yesterday, that this wasn't the correct way to lock the loop as box now expected the gain settings for the loop to be set remotely (and we were providing none of that to the unit). Still, the default gain in REMOTE was enough to provide a stable lock and we didn't understand exactly how that box worked (which is obvious in retrospect). So, yesterday, we pored over the schematics for the TTFSS boxes (Rich is drawing a very nice block diagram to show the loop structure), and realized our error. The correct way to lock is the following:

  • Set to LOCAL & TEST
  • Adjust the temperature so we were close to TEM00 resonance.
  • Scan through the two polarization TEM00 eigenmodes (separated by ~1MHz)
  • We can sit (a) outside the mode with big transmission (big mode), (b) in between the two modes or (c) outside the small mode. 
  • We sit outside the big mode and then switch TEST to OFF to turn the loop on (I have been cursing about the naming conventions on these boxes for the last two days).
  • This locks the loop.
  • With the gains set locally.

From here, we were able to play with the common mode gain settings reduce the noise of the beat note between the lasers. And we were able to lock the PLL. The main evidence for the latter is the fact that we can change the DC value of the control signal in the PLL by varying the carrier frequency of the Marconi.

  1569   Tue Aug 11 16:49:55 2015 Aidan, Rich, AntonioSummaryBEATDiagnosing the terrible beat signal. RCAV HV supply not showing current drawn

The beat signal looks awful. It has some amplitude modulation at 6.75MHz and looks like it has some strange saturation effects going on. This is too much noise for the PLL to lock to.

We thought, for a minute, that the reason for this may be related to one of the HV supplies for the RCAV locking. The needles on the front of the positive supply unit +150V and 0mA current drawn. The other 3 HV supplies in use all show around 20-25mA current draw when used with the TTFSS boards.

We popped the top of the RCAV TTFSS box on the table and looked at the TP4 output on the HV/Interface board (this looks at the signal coming out of the high voltage amplifier that feeds the EOM, but reduced a voltage divider to 1/10th the value). It was freely swinging between +/-4V, so the HV amplifier seems to be happily getting both +ve and -ve voltages. There might be a problem with the needle on the HV supply.

 

Attachment 1: photo_1.JPG
photo_1.JPG
Attachment 2: photo_2.JPG
photo_2.JPG
  1568   Mon Aug 10 18:44:01 2015 Antonio, AidanSummaryFSSFSS interface board

We figured out that when we sweep the laser PZT and we want to look at the error-signal, this can be monitored at the "mixer" output of the FSS interface board while the switch test/off/ramp is on ramp. However we need to understand what this switch does when it is OFF.

 

 

 

 

 

 

Attachment 2: D0902049-v1(front_panel).pdf
D0902049-v1(front_panel).pdf
  1567   Mon Aug 10 17:49:41 2015 AntonioSummaryEOMerror-signal end EOM

On saturday a qualitative effect of the modulation produced by the EOM located in the PDH-north loop has been checked.

The goal was to have a look at the error signal of the PDH-north while the laser PZT was scanning frequecies around the two s-p TEM00 resonances. Because a that time I did not find the right error-signal connections on the FSS board (next elog will clarify where it is) I have demodulated the signal with an external mixer (and with a low pass filter) and monitored it. The picture shows the error-signal that we have with this setup: 

Caption

 

 

 

 

 

  1566   Thu Aug 6 23:49:10 2015 AntonioSummaryFSSpower transmitted through the cavities and control signal

The light transmitted from both the cavities has been monitored while the cavities where locked (Vacav = 1.369 V, Vrcav = 5.7909 V) and beats on RF photodiodes where visible. Power on the three photodiodes PD-Rcav (North cavity), PD-Acav (South cavity) and PD-RF decreases of about 10% from its maximum value on PD-RF and  about 5% on the others photodiodes periodically every ~7 minutes. 

I also notice that:

  1. The PD-Rcav trace is noisier than PD-Acav trace;
  2. The mean voltage values of the two photodiodes are way different:

PD-Rcav = ~60mV (+- 2/3%);

PD-Acav = ~170mV (+-2/3%);

       3. Enabling/disabling the boost switch on the FSS box does not give any improvement;

       4. Pressing the red botton (gain) on the FSS box neither;

 

In the same condition of locking the control signals of both PDH loop have been monitored too.

Here we can see that:

  1. the PDH-North loop is noisier;
  2. A step occurred in the cian trace (at the second division) without causing variation in the transmitted light;
  3. Enabling/disabling the boost switch on the FSS box does not give any improvement;
  4. Pressing the red botton (gain) on the FSS box neither aside from the fact that when the botton is kept pushed it slightly "cleans" the trace (not quantified);

It is worth also monitoring the error signal while the cavities are locked and when they are not (with a triangular wave applied at the laser PZT).

 

Control signal of PDH-North (yellow); Control signal of PDH-South (cian); PD-RF power (pink)
PD_RF
PD-RF
PD-Rcav (lower) , PD-Acav (upper)

 

Attachment 2: Windows_Phone_20150806_18_30_46_Pro.jpg
Windows_Phone_20150806_18_30_46_Pro.jpg
  1565   Wed Aug 5 15:59:07 2015 Aidan, AntonioSummaryRefCavACAV transmission on RF photodiode

We swept the temperature of the ACAV laser and monitored the transmission through the cavity. In the attached image, the TEM00 modes (separated by 3GHz or 1 FSR) are located at C3:PSL-ACAV_SLOWOUT values of 0.5344V and 1.379V. There is a TEM20 mode at 1.52V.


 

Attachment 1: ACAV_FSr_full.png
ACAV_FSr_full.png
  1564   Wed Aug 5 10:18:54 2015 AidanSummaryComputersSummary of installation of Acromag units and control workstation

Summary of installation of new hardware:

 

  1. ACROMAG1 computer.
    1. This is a desktop running a new install of Ubuntu 12.
    2. It is on the CTN lab network and has been assigned the static IP address 10.0.0.33, with a Gateway 255.255.255.0
    3. We can’t seem to access the regular internet on this machine just yet.
    4. I’ve set up the standard CONTROLS account
    5. I’ve installed EPICS per https://dcc.ligo.org/LIGO-T1400200 that includes the MODBUS app that can access the ACROMAG units.
  2. ACROMAG units
    1. The XT1221 (ADC) and XT1541 (DAC) units are installed in the northwest rack
    2. They have static IP addresses:

i.XT1221: 10.0.0.42

ii.XT1541: 10.0.0.41

  1. They are getting a temporary +18V to power them.
  2. The XT1221 is attached to the router and the XT1541 is daisy-chained to the former.
  3. I need to check the averaging settings on the ADC unit to make sure they’re not overtaxing the system
  1. MODBUS/EPICS software
    1. I set up a MODBUS IOC per https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:resources:computing:acromag
    2. The interface to the ACROMAG units is defined in acromag.cmd
    3. The input channels are defined in IOCTEST.DB and these report a value in counts

i.E.g. C3:ACROMAG_INPUT0

  1. We calibrated created three CALC channels that provide outputs calibrated in VOLTS. The calibrations are hardcoded into the database file.

i.C3:PSL-TRANS_ACAV_ISS_DC

ii.C3:PSL-TRANS_RCAV_ISS_DC

iii.C3:PSL-TRANS_RF_DC

  1. The two ISS channels are connected to their respective PDs
  2. The RF_DC output goes through an SR560 with a 30Hz pole and 50x gain on it to amplify the signal before the ADC.
  1563   Tue Aug 4 18:37:21 2015 EvanSummaryRefCavDAQ running with TRANS channels - strange peaks as we scan the laser frequency
Quote:

We now have DC transmission channels in the frames. I'll post the details soon, but here's a plot that shows the transmission through the cavity on the RF_DC channel and the ACAV ISS DC as we slowly ramp the ACAV_SLOWOUT (temperature control) at a rate of 5E-5V/s.

Note the strange shape of the transmission peaks.

The wide view allows us to see the cavity transmit the upper and lower sidebands. Look carefully at the TRANS_RF_DC curve.

 
Bizarro.
 
Do you see the same thing if you unhook the PDs from the digital system and instead watch them on a scope?
 
Do you see the same thing if, instead of sweeping the laser temperature, you put a triangle wave on the laser PZT and scan across the resonances? My recollection is that the resonances (viewed on a scope) should be quite clean and lorentzian, apart from the issue of the two polarizations.
  1562   Tue Aug 4 17:41:01 2015 Aidan, AntonioSummaryRefCavDAQ running with TRANS channels - strange peaks as we scan the laser frequency

We now have DC transmission channels in the frames. I'll post the details soon, but here's a plot that shows the transmission through the cavity on the RF_DC channel and the ACAV ISS DC as we slowly ramp the ACAV_SLOWOUT (temperature control) at a rate of 5E-5V/s.

Note the strange shape of the transmission peaks.

The wide view allows us to see the cavity transmit the upper and lower sidebands. Look carefully at the TRANS_RF_DC curve.

Attachment 1: ACAV_TRANS.png
ACAV_TRANS.png
Attachment 2: ACAV_TRANS_wide.png
ACAV_TRANS_wide.png
  1561   Tue Aug 4 14:15:14 2015 AidanSummaryComputersDAQD not running on FB2. FB2 trend disk is full

Like the title says ...

If I try running DAQD per https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:experiments:psl:add_channel_for_daq_in_fb2

then it fails and the log file reveals that this is when it tries to write a GWF file to the trend folder. Manually trying to write anything to this location results in a "disk full" message.

The trusty df command yields the following.

[controls@fb2 frames]$ df
Filesystem           1K-blocks      Used Available Use% Mounted on
/dev/mapper/VolGroup00-LogVol00
                     232477448 125679200  94798608  58% /
/dev/sda1               101086     25961     69906  28% /boot
tmpfs                  1029664         0   1029664   0% /dev/shm
/dev/sdc1            240362656 205728852  34633804  86% /frames/full
/dev/sdd1            307663800 307663800         0 100% /frames/trend

 

 

  1560   Tue Aug 4 11:28:48 2015 Antonio, AidanSummaryEnvironmentDC transmission channels to Acromag and floor cleaning

Yesterday the DC transmission (ISS photodiode, North cavity) channel to acromag has been added and the viewer has been set on the Ubuntu machine. The channel has been also calibrated (against voltage externally injected). We are going to add other channels from the VME crates and work on the PDH board interface in order to remote control this unit.

Additionally the floor has been sweeped and mopped.

 

 

 

 

 

 

  1559   Sat Aug 1 09:29:25 2015 AidanSummaryEnvironmentCleaned computer table. Added Acromag

I cleaned the table with the computers and removed one of the monitors. I installed the Acromag units to the rack, powered them up and got them onto the network. They are:

XT1221 unit: 10.0.0.42

XT1521 unit: 10.0.0.41

I've pinged them successfully from other machines. I have an Ubuntu machine that I borrowed from the TCS lab to interface to them. I'll set up the EPICS control on Monday. In addition to adding the DC transmission channels to Acromag, we should be able to start migrating the PID controls away from the VME crates to these new units. 

 

 

  1558   Sat Aug 1 09:24:58 2015 Aidan, AntonioSummaryBEATLasers locked to cavities - no beat - polarization? Transmission issues

Both lasers have been locked to the cavities for 24 hours. The slow control of the frequency is handed off to the PID loop. Antonio and I observed strange behaviour on the DC value of the cavity transmission.

As Evan had noted before, there are two polarizations that will resonate and they're about 3MHz apart (if I remember correctly). We can see these on the DC photodiodes on transmission (the ISS PD and the RF DC output). One peak is large and the other much smaller. However, when we have large transmission onto the RF photodiode we have small transmission onto the ISS PD and vice versa. It's likely we have a pick-off optic with strong polarization selectivity.

We couldn't find the beat yesterday or Thursday.

ELOG V3.1.3-